Round the solution to
two decimal places. Can you show me step by step how to do these problems please?
1) 4x – 5 = 16
2) 35 + 3x – 11 = 23

Answer by mananth(16946)

you have to keep the variables on one side of the equation and the numbers/constants on other side

4x – 5 = 16
Adding 5 to both sides eliminates 5 on the left side of the equation
4x-5 +5=16+5
4x = 21
dividing by 4 on both sides we get
1x = 21/4
x= 5.24


35 + 3x – 11 = 23
24+3x =23
add -24 to both sides
24-24+3x=23-24
3x=-1
divide by 3
x=-1/3